OpenStudy (anonymous):

Please help me with this question : Solve using the relate to a simpler problem stradegy . The population of pincrest is 18,000 Lakewoods population 21/25 of pintrest What is lakewoods population? A. 21,149 B. 15,120 C. 741 D. 721

OpenStudy (anonymous):

@amistre64 @abb0t @Ashleyisakitty @bibby @chucky00 @J-bird

OpenStudy (amistre64):

um, just multiply the information ...

OpenStudy (anonymous):

I got A is that right

OpenStudy (amistre64):

no, if something is less than 100% of something else, that it cannot be more than 100% of it

OpenStudy (amistre64):

21/25 = 21(4)/100 = 84%

OpenStudy (anonymous):

basically 21/25=84%. to multiply a percent we put into a decimal. 0.84 of 18000

OpenStudy (amistre64):

10% is 1800 so we are somewhere between 1800 and 18000

OpenStudy (amistre64):

if we divide 18000 into 25 parts 18000/25 and add up 21 of those parts 18000 * 21/25

OpenStudy (anonymous):

b then

OpenStudy (amistre64):

b is ht emost reasonable to me yes